### v2

@@ C.1 Auxiliary lemmas
The proof completes by recognizing that the $o(1/K_{n})$ term is dominated by the first term.
C.6 Proof of Theorem 3.4
We prove the following more general result:
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}+\lambda\Big{(}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}=\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)(1+o(1)).$
Clearly, the left-hand side is upper bounded by $\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)$ which, by Theorem 3.3, is equal to the right-hand side. From the proofs of Lemma 3.1 and Lemma C.2, we observe that the worst-case bias and variance are both achieved, up to $1+o(1)$ factors, by the potential outcomes defined in (28). The result is then proved. In particular, when $\lambda=1$, the left-hand side is the worst-case MSE because
$\mathbb{E}_{\mu}[(\hat{\tau}_{n}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]=\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}+\Big{(}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}.$
C.7 Proof of Theorem 3.5
We decompose the proof as follows. First, we characterize the bias and variance of the Bernoulli design. We then compare $\mathcal{B}_{n}$ for the cluster and Bernoulli design, assuming that $\lambda\in(0,\infty)$ is bounded away from zero and infinity as assumed in Theorem 3.5.
Worst-case bias and variance of the Bernoulli and cluster design
Note that $\mathcal{C}_{\mathrm{B},n}$ satisfies Assumption 4. By Lemma C.1 and C.2, $\mathcal{C}_{\mathrm{B},n}$ has worst-case bias $\bar{\phi}_{n}b_{n}(\mathcal{C}_{\mathrm{B},n})=\bar{\phi}_{n}$, and worst-case variance $\frac{1}{n}[\bar{\psi}+o(1)]$. Thus,
@@
Case $\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)>\mathcal{B}_{n}(\mathcal{C}_{\mathrm{B},n},\lambda)$.
By Theorem 3.3, it remains to prove that
$\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)\geq\frac{\bar{\psi}}{K_{n}},$
and
Case $\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)<\mathcal{B}_{n}(\mathcal{C}_{\mathrm{B},n},\lambda)$
Similarly, by Theorem 3.3, it remains to prove that
and
C.8 Proof of Theorem 3.6
The proof follows directly the proof of Theorem 3.5 (case for $\mathcal{B}_{n}(\mathcal{C}_{\mathrm{B},n},\lambda)\geq\mathcal{B}_{n}(\mathcal{C}_{n},\lambda)$), after noticing that the result directly holds if we replace $\bar{\gamma}^{2}$ in the proof of Theorem 3.5 with $\underline{\gamma}$ as defined in the statement of Theorem 3.6, and rearranging terms accordingly.
C.9 Proof of Theorem 4.1
@@
$\min_{c:\{1,\cdots,n\}\mapsto\{1,\cdots,K\}}-\text{tr}(\mathbf{M}_{c}^{T}n\mathbf{L}\mathbf{M}_{c})+\xi_{n}(\mathbf{1}_{n}^{T}\mathbf{M}_{c}\mathbf{M}_{c}^{T}\mathbf{1}_{n})$
$=\min_{c:\{1,\cdots,n\}\mapsto\{1,\cdots,K\}}-\text{tr}((n\mathbf{L}-\xi_{n}\mathbf{1}_{n}\mathbf{1}_{n}^{T})\mathbf{M}_{c}\mathbf{M}_{c}^{T}).$
C.10 Proof of Theorem 6.1
We study the worst-case bias and variance separately. We conclude with a characterization of the MSE.
@@
Again, the second term is the variance of the difference in means estimator with known $\mu_{i}(\mathbf{d})$’s. By (36) in the proof of Lemma C.2,
@@
Following the proof of Lemma C.2, we can recover (36):
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}+o(1/K_{n}).$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\leq\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}+o(1/K_{n}).$
@@
where the last line applies Assumption 4 for the second term and Cauchy-Schwarz inequality for the third term. Since $(1/n)\sum_{i=1}^{n}\bar{\psi}_{i}\leq\bar{\psi}=O(1)$ and $\bar{\phi}_{n}=o(1)$ by Assumption 3,
$\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}\geq\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}+o(1/K_{n})$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}+o(1/K_{n}).$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}(1+o(1)).$
@@
Thus, because $\bar{\phi}_{n}=o(1)$ under Assumption 3, $\bar{\phi}_{n}\sum_{i=1}^{n}\Big{|}\frac{w_{i}}{\|w\|_{1}}-\frac{1}{n}\Big{|}=o(U_{n}^{*}(\mathcal{C}_{n},w))$ and
$|\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}]-\tau_{n,\mu}|=U_{n}^{*}(\mathcal{C}_{n},w)(1+o(1)).$
@@
Following the proof of Lemma C.2, we obtain that
@@
$\frac{\bar{\phi}_{n}}{n^{2}}\cdot\sum_{i}|B_{i}\cup G_{i}|=o(1/K_{n}).$
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}+o(1/K_{n}).$ (49)
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\leq\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\bar{\psi}+o(1/K_{n}).$
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}$ $=\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot(\bar{\psi}^{1/2}-\bar{\phi}_{n})^{2}+o(1/K_{n})$
$=\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\bar{\psi}(1+o(1))+o(1/K_{n}).$

### v3

@@ D.1 Auxiliary lemmas
The proof completes.
D.6 Proof of Theorem 3.3
Clearly from Lemma 3.1 and 3.2 we have
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}+\lambda\Big{(}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\leq\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)(1+\mathcal{O}(\varepsilon)).$
We now have to show the other side of the inequality. From the proofs of Lemma 3.1 and Lemma D.2, we observe that the worst-case bias and variance are both achieved, up to $1+\mathcal{O}(\varepsilon)$ factors, by the potential outcomes defined in Equation (33). The result is then proved.
D.7 Proof of Theorem 3.4
We can write,
@@
Finally, note that $\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}=(\frac{1}{K_{n}}\sum_{k=1}^{K_{n}}\gamma_{k}^{2})/K_{n}=\underline{\gamma}/K_{n}$. The proof completes after rearrangement.
D.8 Proof of Theorem 4.1
@@
$\min_{c:\{1,\cdots,n\}\mapsto\{1,\cdots,K\}}-\text{tr}(\mathbf{M}_{c}^{T}n\mathbf{L}\mathbf{M}_{c})+\xi(\mathbf{1}_{n}^{T}\mathbf{M}_{c}\mathbf{M}_{c}^{T}\mathbf{1}_{n})$
$=\min_{c:\{1,\cdots,n\}\mapsto\{1,\cdots,K\}}-\text{tr}((n\mathbf{L}-\xi\mathbf{1}_{n}\mathbf{1}_{n}^{T})\mathbf{M}_{c}\mathbf{M}_{c}^{T}).$
Appendix E Proofs for extensions
E.1 Auxiliary lemmas for extensions
Lemma E.1.
Let $X_{1},\ldots,X_{n}$ be a set of random variables whose dependency graph has degree $m_{i}$ for variable $X_{i}$, i.e.,
|$|\{j\in\{1,\ldots,n\}:X_{j}\not\perp\!\!\!\perp X_{i}\}|=m_{i}.$
Let $m_{\max}=\max_{i}m_{i}$ denote the maximal degree. Assume that
$\max_{i}|X_{i}|\leq B,$
for some constant $B$ that does not depend on $n$, and
$\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[X_{i}]\right|\leq b_{n},\quad\mathrm{Var}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\right)\leq\sigma_{n}^{2},$
for some deterministic sequences $b_{n},\sigma_{n}$ that depend on $n$. Further assume that, as $n\rightarrow\infty$,
$\frac{1}{\sigma_{n}^{3}}\frac{m_{\max}^{2}}{n^{2}}+\frac{1}{\sigma_{n}^{2}}\frac{m_{\max}^{3/2}}{n^{3/2}}=o(1).$ (43)
Then
$\liminf_{n\rightarrow\infty}\mathbb{P}\left(\left|\frac{1}{n}\sum_{i=1}^{n}X_{n}\right|\leq\chi(b_{n},\sigma_{n})\right)\geq 1-\alpha.$
Proof.
Let
$S_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i},\quad\nu_{n}^{2}=\mathrm{Var}(S_{n}),\quad\mu_{n}=\mathbb{E}[S_{n}].$
First, note that from Theorem 3.5 in Ross (2011),
$d(S_{n},N(\mu_{n},\nu_{n}^{2}))=O\left(\frac{m_{\max}^{2}}{\nu_{n}^{3}n^{2}}+\frac{m_{\max}^{3/2}}{\nu_{n}^{2}n^{3/2}}\right),$
where $d(\cdot)$ is the Wasserstein distance.
Thus, because the Wasserstain distance is an upper bound to the Kolmogorv distance for two random variables $(X,Y)$ provided $Y$ is Gaussian (e.g. Gaunt and Li, 2023), we can write
$\mathbb{P}\left(|S_{n}|>\chi(|\mu_{n}|,\nu_{n})\right)$ $\leq\mathbb{P}\left(|N(\mu_{n},\nu_{n}^{2})|>\chi(|\mu_{n}|,\nu_{n})\right)+O\left(\frac{m_{\max}^{2}}{\nu_{n}^{3}n^{2}}+\frac{m_{\max}^{3/2}}{\nu_{n}^{2}n^{3/2}}\right)$
$=\alpha+O\left(\frac{m_{\max}^{2}}{\nu_{n}^{3}n^{2}}+\frac{m_{\max}^{3/2}}{\nu_{n}^{2}n^{3/2}}\right).$
By Lemma E.3,
$\mathbb{P}\left(|S_{n}|>\chi(b_{n},\sigma_{n})\right)\leq\alpha+O\left(\frac{m_{\max}^{2}}{\nu_{n}^{3}n^{2}}+\frac{m_{\max}^{3/2}}{\nu_{n}^{2}n^{3/2}}\right).$ (44)
On the other hand, by triangle inquality
$\mathbb{P}\left(|S_{n}|>\chi(b_{n},\sigma_{n})\right)\leq\mathbb{P}\left(|S_{n}-\mu_{n}|>\chi(b_{n},\sigma_{n})-|\mu_{n}|\right)\leq\mathbb{P}\left(|S_{n}-\mu_{n}|>\chi(b_{n},\sigma_{n})-b_{n}\right).$
By Lemma E.2,
$\chi(b_{n},\sigma_{n})-b_{n}\geq q_{1-\alpha}\sigma_{n},$
where $q_{1-\alpha}$ is the $1-\alpha$ quantile of a standard Normal random variable. By Cherbyshev’s inequality,
$\mathbb{P}\left(|S_{n}|>\chi(b_{n},\sigma_{n})\right)\leq\mathbb{P}\left(|S_{n}-\mu_{n}|>q_{1-\alpha}\sigma_{n}\right)\leq\frac{\nu_{n}^{2}}{q_{1-\alpha}^{2}\sigma_{n}^{2}}=O\left(\frac{\nu_{n}^{2}}{\sigma_{n}^{2}}\right).$ (45)
Fix any deterministic sequence $a_{n}$. If $\nu_{n}<a_{n}\sigma_{n}$, (45) implies
$\mathbb{P}\left(|S_{n}|>\chi(b_{n},\sigma_{n})\right)=O(a_{n}^{2}).$
If $\nu_{n}\geq a_{n}\sigma_{n}$, (44) implies
$\mathbb{P}\left(|S_{n}|>\chi(b_{n},\sigma_{n})\right)\leq\alpha+O\left(\frac{1}{a_{n}^{3}}\Big{(}\frac{m_{\max}^{2}}{\sigma_{n}^{3}n^{2}}+\frac{m_{\max}^{3/2}}{\sigma_{n}^{2}n^{3/2}}\Big{)}\right).$
Define $c_{n}=\frac{nm_{\max}^{2}}{\sigma_{n}^{3}}+\frac{n^{1/2}m_{\max}^{3/2}}{\sigma_{n}^{2}}$. Equation (43) implies $c_{n}=o(1)$. Setting $a_{n}=c_{n}^{1/4}$ yields
$\mathbb{P}\left(|S_{n}|>\chi(b_{n},\sigma_{n})\right)\leq\alpha+O\left(\min\left\{a_{n}^{2},\frac{c_{n}}{a_{n}}\right\}\right)=\alpha+o(1).$
E.1.1 Proof of Lemma 6.1
Define
$B_{i}=\{v:\text{either }c(v)=c(i)\text{ or }c(v)=c(v^{\prime})\text{ for some }v^{\prime}\in\mathcal{N}_{i}\},\quad G_{i}=\{g:\mathcal{N}_{g}\cap B_{i}\neq\emptyset\}.$
Lemma D.1 implies
$\mu_{j}(\mathbf{D})\perp\!\!\!\perp\mu_{i}(\mathbf{D}),\quad\forall j\not\in B_{i}\cup G_{i}.$ (46)
It is easy to see that
$j\in B_{i}\cup G_{i}\Longleftrightarrow i\in B_{j}\cup G_{j}.$
Thus, (46) defines a undirected dependency graph among $(\mu_{1}(\mathbf{D}),\ldots,\mu_{n}(\mathbf{D}))$. Define
$i\sim j\text{ iff }j\in B_{i}\cup G_{i}.$
Under our neighborhood interference assumption, we have shown in the proof of Lemma D.1 that each summand in $\hat{\tau}$ forms a local dependency graph with degree $m_{i}$ as defined in Lemma E.1 such that
$m_{i}$ $\leq\sum_{\ell:\ell\in B_{i}}(|\mathcal{N}_{\ell}|+1)\leq|B_{i}|\cdot(\mathcal{N}_{n,\max}+1)$
$\leq\max_{k}n_{k}(|\mathcal{N}_{i}|+1)(\mathcal{N}_{n,\max}+1)\leq 4\bar{\gamma}\mathcal{N}_{n,\max}^{2}\frac{n}{K_{n}}.$ (47)
In addition, by assumption $\bar{\sigma}_{n}^{2}\geq\frac{\eta+o(1)}{K_{n}}$ for a constant $\eta>0$. We now invoke Lemma E.1. Note that we can write
$\hat{\tau}_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$
where $X_{i}$ are uniformly bounded random variables forming a dependency graph with maximum degree $m_{\max}=\mathcal{O}(\mathcal{N}_{n,\max}^{2}\frac{n}{K_{n}}$). To check that the conditions in Lemma E.1 we only need to show that
$\frac{1}{\bar{\sigma}_{n}^{3}n^{2}}m_{\max}^{2}+\frac{1}{\bar{\sigma}_{n}^{2}n^{3/2}}m_{\max}^{3/2}=\mathcal{O}\left(\frac{K_{n}^{3/2}}{n^{2}}m_{\max}^{2}+\frac{K_{n}}{n^{3/2}}m_{\max}^{3/2}\right)=\mathcal{O}\Big{(}\frac{\mathcal{N}_{n,\max}^{2}}{K_{n}^{1/2}}\Big{)}$
is $o(1)$. This follows by assumption, completing the proof.
E.2 Proof of Theorem 6.2
Before proving the theorem we need to establish some properties.
E.2.1 Useful properties of $\chi(b,\nu)$
Let $\Phi(\cdot)$ denote the CDF of $N(0,1)$. By definition,
$\Phi\left(\frac{\chi(b,\nu)-b}{\nu}\right)-\Phi\left(\frac{-\chi(b,\nu)-b}{\nu}\right)=1-\alpha.$ (48)
Throughout the section we assume $b>0$, which is true in (15).
Lemma E.2.
Assume $\alpha\leq 0.5$. Let $z_{\beta}$ denote the $\beta$-th quantile of $N(0,1)$. Then
$\chi(b,\nu)\in[b+q_{1-\alpha}\nu,b+q_{1-\alpha/2}\nu].$
Proof.
Since $\Phi(0)=0.5\geq\alpha$,
$\chi(b,\nu)\geq b>0.$
Thus,
$\Phi\left(\frac{\chi(b,\nu)-b}{\nu}\right)-\Phi\left(\frac{-\chi(b,\nu)-b}{\nu}\right)\leq 1-\alpha\leq\Phi\left(\frac{\chi(b,\nu)-b}{\nu}\right).$
Using the symmetry of $N(0,1)$, we have
$21-\alpha\leq\Phi\left(\frac{\chi(b,\nu)-b}{\nu}\right)\leq 1-\alpha/$ 2 .
The proof is then completed. ∎
Lemma E.3.
$\chi(b,\nu)$ is increasing in $b$ and $\nu$.
Proof.
Let $\chi_{b}(b,\nu)$ and $\chi_{\nu}(b,\nu)$ denote the partial derivatives of $\chi$ with respect to $b$ and $\nu$, respectively. We suppress the dependence on $(b,\nu)$ in $\chi,\chi_{b},\chi_{\nu}$ for notational convenience. Let $\phi(x)=\Phi^{\prime}(x)=1/\sqrt{2\pi}\exp\{-x^{2}/2\}$. Differentiating (48) with respect to $b$, we obtain that
$0=(\chi_{b}-1)\phi\left(\frac{\chi-b}{\nu}\right)+(\chi_{b}+1)\phi\left(-\frac{\chi+b}{\nu}\right)=(\chi_{b}-1)\phi\left(\frac{\chi-b}{\nu}\right)+(\chi_{b}+1)\phi\left(\frac{\chi+b}{\nu}\right),$
where the last line uses the symmetry of $\phi$. As a result,
$\chi_{b}=\frac{\phi\left(\frac{\chi-b}{\nu}\right)-\phi\left(\frac{\chi+b}{\nu}\right)}{\phi\left(\frac{\chi-b}{\nu}\right)+\phi\left(\frac{\chi+b}{\nu}\right)}.$ (49)
Since $\chi\geq b>0$ and $\phi$ is decreasing on $[0,\infty)$,
$\phi\left(\frac{\chi-b}{\nu}\right)>\phi\left(\frac{\chi+b}{\nu}\right).$
Thus, $\chi_{b}>0$, implying that $\chi$ is increasing in $b$.
Differentiating (48) with respect to $\nu$, we obtain that
00 0 $=\frac{\nu\chi_{\nu}-\chi+b}{\nu^{2}}\phi\left(\frac{\chi-b}{\nu}\right)+\frac{\nu\chi_{\nu}-\chi-b}{\nu^{2}}\phi\left(-\frac{\chi+b}{\nu}\right)$
$=\frac{\nu\chi_{\nu}-\chi+b}{\nu^{2}}\phi\left(\frac{\chi-b}{\nu}\right)+\frac{\nu\chi_{\nu}-\chi-b}{\nu^{2}}\phi\left(\frac{\chi+b}{\nu}\right),$
where, again, the last line uses the symmetry of $\phi$. Tis implies
$\chi_{\nu}=\frac{\frac{\chi-b}{\nu}\phi\left(\frac{\chi-b}{\nu}\right)+\frac{\chi+b}{\nu}\phi\left(\frac{\chi+b}{\nu}\right)}{\phi\left(\frac{\chi-b}{\nu}\right)+\phi\left(\frac{\chi+b}{\nu}\right)}>0.$ (50)
Thus, $\chi$ is increasing in $\nu$. ∎
Lemma E.4.
$\chi(b,\nu)$ is Lipschitz in $b$ and $\nu$. The Lipschitz constant is bounded by $L_{\alpha}$ for some constant $L_{\alpha}$ that only depends on $\alpha$.
Proof.
By (49), it is clear that
$\chi_{b}(b,\nu)\leq 1,\quad\forall b,\nu\geq 0.$
Let $a=(\chi(b,\nu)-b)/\nu,w=2b/\nu$. Then (50) can be rewritten as
$\chi_{\nu}(b,\nu)=\frac{a\phi(a)+(a+z)\phi(a+z)}{\phi(a)+\phi(a+z)}=a+\frac{z}{1+\phi(a)/\phi(a+z)}=a+\frac{z}{1+\exp\{az+z^{2}/2\}}.$
Using the inequality that $\exp\{y\}\geq 1+y$, we have
$\chi_{\nu}(b,\nu)\leq a+\frac{z}{2+z^{2}/2+az}\leq a+\frac{1}{a+2}\leq a+0.5.$
By Lemma E.2, we conclude that
$\chi_{\nu}(b,\nu)\leq q_{1-\alpha/2}+0.5,\quad\forall b,\nu\geq 0.$
By Lemma E.3, $\chi_{b}(b,\nu),\chi_{\nu}(b,\nu)\geq 0$. Thus, the Lipschitz constant is bounded by
$\sqrt{\chi_{b}(b,\nu)^{2}+\chi_{\nu}^{2}}\leq\sqrt{1+(q_{1-\alpha/2}+0.5)^{2}}.$
E.2.2 Proof of the theorem
By Lemma E.3,
$|I_{\mathrm{ea}}(\beta(\mu),\nu(\mu))|\leq|I_{\mathrm{ea}}^{*}|.$ (51)
In the proof of worst-case bias and variance, we have shown there exists $\mu^{*}$ such that
$|\beta(\mu^{*})|=b_{n,*},\quad\sigma^{2}(\mu^{*})=\sigma_{n,*}^{2}(1+\mathcal{O}(\varepsilon)).$
By Lemma E.4, from a first order Taylor approximation so that $\sigma_{n,*}\sqrt{(1+\mathcal{O}(\varepsilon))}=\sigma_{n,*}+\mathcal{O}(\varepsilon)$
$|I(\mu^{*})|\geq|I_{\mathrm{ea}}^{*}|\cdot(1+\mathcal{O}(\varepsilon)).$ (52)
Combining (51) and (52), we obtain the following result.
E.3 Proof of Theorem B.1
We study the worst-case bias and variance separately. We conclude with a characterization of the MSE. Recall that for simplicity here we take a local asymptotic framework with $\varepsilon=0$, although our derivations can be directly extended to positive $\varepsilon$ as for the other proofs.
@@ E.2.2 Proof of the theorem
Again, the second term is the variance of the difference in means estimator with known $\mu_{i}(\mathbf{d})$’s. By (41) in the proof of Lemma D.2,
@@
Following the proof of Lemma D.2, we can recover (41):
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}+\mathcal{O}(\varepsilon/K_{n}).$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\leq\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}+\mathcal{O}(\varepsilon/K_{n}).$
@@
where the last line applies Assumption 4 for the second term and Cauchy-Schwarz inequality for the third term. Since $(1/n)\sum_{i=1}^{n}\bar{\psi}_{i}\leq\bar{\psi}=O(1)$ and $\bar{\phi}_{n}=\mathcal{O}(\varepsilon)$ by Assumption 3,
$\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}\geq\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}+\mathcal{O}(\varepsilon/K_{n})$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}+\mathcal{O}(\varepsilon/K_{n}).$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}(1+\mathcal{O}(\varepsilon)).$
@@
Thus, because $\bar{\phi}_{n}=\mathcal{O}(\varepsilon)$ under Assumption 3, $\bar{\phi}_{n}\sum_{i=1}^{n}\Big{|}\frac{w_{i}}{\|w\|_{1}}-\frac{1}{n}\Big{|}=\mathcal{O}(\varepsilon U_{n}^{*}(\mathcal{C}_{n},w))$ and
$|\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}]-\tau_{n,\mu}|=U_{n}^{*}(\mathcal{C}_{n},w)(1+\mathcal{O}(\varepsilon)).$
@@
Following the proof of Lemma D.2, we obtain that
@@
$\frac{\bar{\phi}_{n}}{n^{2}}\cdot\sum_{i}|B_{i}\cup G_{i}|=\mathcal{O}(\varepsilon/K_{n}).$
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}+\mathcal{O}(\varepsilon/K_{n}).$ (64)
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\leq\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\bar{\psi}+\mathcal{O}(\varepsilon/K_{n}).$
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}$ $=\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot(\bar{\psi}^{1/2}-\bar{\phi}_{n})^{2}+\mathcal{O}(\varepsilon/K_{n})$
$=\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\bar{\psi}(1+o(1))+\mathcal{O}(\varepsilon/K_{n}).$